OpenStudy (anonymous):

multiply and simplify x^2-3x-28/x^2-8x+16 * x-4/x-7 =

OpenStudy (anonymous):

Hey Jazil, this is similar to another question you asked. I think what you might like is to keep these few things in mind: 1.) Factor if able - See if you can factor out things where x is to a power like \[x^2\] 2.) Whenever you divide by a fraction, you can multiply by the reciprocal **Not needed for this problem, but useful for the problems you're asking** So try this question on your own with this knowledge and then look down below to compare it with my work: \[\frac{(x^2-3x-28)}{(x^2-8x+16)} * \frac{(x-4)}{(x-7)}\] \[\frac{(x-7)(x+4)}{(x-4)(x-4)} * \frac{(x-4)}{(x-7)}\] Things cancel out, this is the fun part. \[\frac{(x+4)}{(x-4)}\] Pretty fancy?

OpenStudy (anonymous):

i came up with x^2 - 28/x^2 + 5x - 4/x - 7

OpenStudy (anonymous):

Hmm, I can't follow your work there. It's important that you don't cancel out things that you can't... It's hard to explain that in writing but think of the part on top of the faction as a whole. If there is plus or minus signs you cannot cancel out things. Only when there is multiplication/division can you cancel. Perhaps type out your steps so that I can see where you went wrong. You have to treat the \[x^2-3x-28\] as one whole thing. It's -3x cannot cancel out with the -8x on the bottom. An easy way to keep track of yourself is only cancel out when you have multiplication or division signs. Do not cancel when there are plus or minus signs separating the variables.